Earleton Manufacturing Company has $3 billion in sales and $800,000,000 in fixed assets. Currently, the company's fixed assets are operating at 80% of capacity. What level of sales could Earleton have obtained if it had been operating at full capacity? Write out your answer completely. Round your answer to the nearest whole number. $ What is Earleton's target fixed assets/sales ratio? Round your answer to two decimal places. % If Earleton's sales increase 35%, how large of an increase in fixed assets will the company need to meet its target fixed assets/sales ratio? Write out your answer completely. Do not round intermediate calculations. Round your answer to the nearest whole number
Answer #1

Level of sales at full capacity = 3,000,000,000/80% = $3,750,000,000

Target ratio = 3,750,000,000/800,000,000

= 4.6875

New sales = 3,000,000,000*1.35 = $4,050,000,000

Fixed assets required = 4,050,000,000/4.6875 = $864,000,000

Hence, increase in fixed assets required = $64,000,000